OpenStudy (anonymous):

Find all solutions to 3x^2 + 2x + 1 = 0

OpenStudy (anonymous):

Please help.

OpenStudy (anonymous):

According to discriminant b^2-4ac. (2)^2 - 4*(3)*1 = -8 < 0 therefore no solutions. See for yourself on the graph. It never touches 0.

OpenStudy (anonymous):

Thank you SO much:) can you help me with some more?

OpenStudy (anonymous):

Sure, bring it on

OpenStudy (anonymous):

Charlie uses the function f(x) = 2x^2 + 5x + 3 to track the temperature of a substance, where x represents the time. When you evaluate f(5) you find that the temperature of the substance is 78°C. What is the domain of the situation of Charlie's quadratic function?

OpenStudy (anonymous):

The function has real values for any x value, so the domain is all real x.

OpenStudy (anonymous):

However, you might want to argue that since the time can never be negative, the domain is x ≥ 0

OpenStudy (anonymous):

Thank you, I wish I was good in math..Write the quadratic function in vertex form. f(x) = x^2 - 3x - 10

OpenStudy (anonymous):

Hmmm, never come across vertex form. Will look it up

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

OK. Put all the numbers (with no x's) on one side with the y x^2 - 3x = y + 10 Start off by completing the square. (what value of ? will make the LHS a perfect square) x^2 - 3x + ? = y + 10 + ? ? = (3/2)^2 = 9/4 x^2 - 3x + 9/4 = y + 10 + 9/4 (x - 1.5)^2 = y + 12.25 And I believe this is now in vertex form

OpenStudy (anonymous):

Or you could do: y = (x - 1.5)^2 - 12.25

OpenStudy (anonymous):

Thank you! The parabola with the equation y = x^2 was both horizontally and vertically translated to produce the parabola with the equation y = x^2 – 14x + 39. Which statement is true? A) The parabola with the equation y = x^2 was translated 7 units to the left and 10 units down. B) The parabola with the equation y = x^2 was translated 7 units to the left and 10 units up. C) The parabola with the equation y = x^2 was translated 7 units to the right and 10 units down. D) The parabola with the equation y = x^2 was translated 7 units to the right and 10 units up.

OpenStudy (anonymous):

ahh. Step 1 FACTORISE

OpenStudy (anonymous):

No, actually not yet

OpenStudy (anonymous):

so dont factor

OpenStudy (anonymous):

I'm thinking...

OpenStudy (anonymous):

x^2 – 14x + 39 = (x^2 - 14x + 49) + 39 = (x-7)^2 - 10 so then we have y = (x-7)^2 - 10 rearrange: y + 10 = (x - 7)^2 Now, the negatives of the numbers are the answers: it moved the graph 10 down the y-axis and 7 right across the x-axis If its a quadratic, you want to get it into (x+?)^2 form.